OpenStudy (mindblast3r):

which store has the lower price for milk, $4.75/gal. or $2.99/2 L carton??

OpenStudy (kkutie7):

So 4.75 for 1 gal and 2.99 for 2L?

OpenStudy (mindblast3r):

Yah.

OpenStudy (kkutie7):

3.785L/gal 4.75/3.785L or. 2.99/2L. Divide both and see what you get.

OpenStudy (mindblast3r):

$1.25 for the first one, $1.50 for the second one.

OpenStudy (mindblast3r):

but that's wrong.. i think lol.

OpenStudy (kkutie7):

Yeah.... Give me a min haha

OpenStudy (kkutie7):

Well another way you can look at it is by picking a number of L bigger than 3 I'd say 5 and see how they compare.

OpenStudy (kkutie7):

5L*($4.75/3.785L)=$ 5L*($2.99/2L)=$

OpenStudy (mindblast3r):

err.. it's going to be store A always with the lowest price lol.

OpenStudy (kkutie7):

Yupp. Does this make sense to you? It barely makes sense to me haha.

OpenStudy (mindblast3r):

when you do 2.99/2 = 1.25 what is this?? 1.25 represents what?

OpenStudy (mindblast3r):

kinda...

OpenStudy (kkutie7):

Dollars per liters

OpenStudy (mindblast3r):

but like 4.75 / 3.79 = $1.50

OpenStudy (kkutie7):

It's like this $4.75/3.79L right

OpenStudy (mindblast3r):

yea

OpenStudy (mindblast3r):

when you divide it gets 1.50$

OpenStudy (mindblast3r):

what does that mean??

OpenStudy (kkutie7):

So when you divide and get 1.50 the units stay so $1.50/L the division sign means per so there are $1.50 per one Liter

OpenStudy (mindblast3r):

are you 100% sure?

OpenStudy (kkutie7):

Yeah I'm pretty sure

OpenStudy (mindblast3r):

alright thanks for helping me. :)

OpenStudy (kkutie7):

No problem xD